# Tight contact structures without symplectic fillings are everywhere

### 27 Feb 2024 v4

@@ 1. Introduction
In contact topology, there is an important dichotomy dividing contact structures into those that are overtwisted and those that are tight. According to [Eli89, BEM15], the former are in fact topological and not geometric objects, in the sense that they satisfy $h$-principles saying that their existence and classification can be reduced to classical obstruction theory and algebraic topology. On the other hand, tight contact structures are inherently geometric, hence in general much harder to construct, and let alone classify. There is however an important sufficient condition for tightness: every contact manifold which is symplectically fillable [^1] Here we consider strong fillings, i.e. symplectic manifolds such that the contact structure at the boundary is defined by the contraction of the symplectic form with a Liouville vector field, locally defined near the boundary. is necessarily tight [Gro85, Eli90, Nie06, BEM15]. A natural question is then to understand the difference between the classes of tight and fillable contact structures. Our first result addresses this question in broad generality, on manifolds of dimension at least five.
@@
If $(M^{2n+1},\xi)$ with $n\geq 3$ is tight, then $M$ admits a tight, non-fillable contact structure in the same almost contact class. If $n=2$, the same holds, with the extra assumption that the first Chern class of $\xi$ is zero.
The assumption on the dimension in Theorem A is necessary, since in dimension three the statement is false. For example the sphere $\mathbb{S}^{3}$ admits a unique tight contact structure [Eli92], which is also symplectically fillable (the standard one). Historically the first examples of tight contact structures that are not symplectically fillable were found in dimension $3$ by Entyre-Honda [EH02] and on a special class of product manifolds by Massot-Niederkrüger-Wendl in dimensions 5 and greater [MNW13].
@@
For every $n\geq 2$, the sphere $\mathbb{S}^{2n+1}$ admits a tight, non-fillable contact structure that is homotopically standard.
Here, homotopically standard means homotopic to the standard contact structure among almost contact structures (the algebraic topological object underlying contact structures, in terms of which the $h$-principles in [Eli89, BEM15] are formulated). We then deduce Theorem A by considering connected sums with these non-fillable examples on spheres to preserve the smooth topology but alter the properties of the contact structure. Of course, this requires some significant effort since both the obstruction to fillability as well as tightness do not obviously behave well under connected sum.
@@
The notion of fillability comes with several different nuances, namely one has notions of Weinstein fillability, Liouville fillability and strong fillability. Thus ordered, each notion is stronger than the following one, and all of them in fact imply tightness. One thus has the following natural hierarchy of contact manifolds:
[…]
There has been a considerable amount of work over almost 30 years in studying these inclusions, and they are known to be strict in any odd dimension [Eli96, EH02, Ghi05, Gay06, NW11, Bow12, MNW13, BCS14, Zho21a, BGM22, GNE22]. On the other hand, all the examples known so far of contact structures with exotic fillability properties come from special geometric constructions, all of which again require the presence of non-trivial topology in the underlying smooth manifold. Now, as a consequence of Theorem A and Theorem B, the final inclusion is strict for all contact manifolds, as soon as the set of tight contact structures is non-empty.
@@
In fact, there is also the further notion of weak fillability, which as the name suggests is the weakest possible notion of fillability. Since this in fact agrees with strong fillability on spheres due to [MNW13, Proposition 6], we do not consider weak fillability here. In particular, even though strictly speaking in (the proof of) Theorem B we obstruct strong fillability, this equivalence of strong and weak fillability for spheres implies that those contact spheres in fact are also not weakly fillable.
@@
By taking connected sum with examples on spheres we can further deduce that there are Liouville fillable contact manifolds without Weinstein fillings in the case that the first Chern class is torsion (cf. Theorem E below). Thus, other than addressing the $5$-dimensional case, it remains to determine the existence of strongly but not Liouville fillable contact structure on $\mathbb{S}^{2n+1}$, for all $n\geq 2$.
@@
After one constructs examples of contact structures with exotic fillability properties, it is natural to try to produce infinitely many such examples as well. In this direction, by pushing further the techniques used for the proofs of the previous statements, namely this time taking contact connected sums with some other judiciously chosen Weinstein fillable contact structures on the sphere, we obtain the following.
@@
Let $n\geq 5$, and suppose that $M^{2n+1}$ admits a Weinstein fillable contact structure with torsion first Chern class. Then, $M$ admits infinitely many non-isomorphic, tight and not strongly fillable contact structures in the same almost contact class.
In particular, spheres of dimension at least $11$ admit infinitely many tight non-fillable contact structures which are homotopically standard. One also has a similar result for Liouville but not Weinstein fillable contact structures.
@@
In order to prove Theorem B we apply flexible contact surgery to examples that are known not to admit strong symplectic fillings, in order to obtain a (smoothly standard) sphere. Namely, we consider the Bourgeois construction on a sphere $\mathbb{S}^{2n-1}$ (with respect to a suitable open book), which yields a contact structure on $\mathbb{S}^{2n-1}\times\mathbb{T}^{2}$, and then kill the topology coming from the $\mathbb{T}^{2}$-factor via surgery. In doing so one has two major obstacles to deal with: namely one must ensure that tightness is preserved, and also that the obstructions to fillability persist. To ensure the former we use a dynamical criteria on the periodic orbits of suitable contact forms, which is a strong version of asymptotic dynamical convexity (ADC) introduced in [Laz20a, Zho21b], that in turn ensures algebraic tightness.
@@
To find fillability obstructions that are not destroyed via surgery, we exploit ideas that have been used to classify fillings of such ADC-manifolds under stronger assumptions (such as symplectic asphericity, vanishing of $c_{1}$). This yields homological restrictions on such putative fillings, which can then be used themselves to provide obstructions. This approach was used for instance in [BGM22] to obstruct fillability of Bourgeois contact manifolds, where one used moduli spaces of closed spheres that appear after applying a certain capping procedure. We instead use a partial cap to build a strong cobordism to a flexibly fillable manifold, which admits a natural filling whose symplectic cohomology is known to vanish. One can then apply arguments classifying exact fillings of such flexible contact manifold as in [Zho22a], to obtain homological obstructions that can be made to persist under connected sum, thus also yielding Theorem A. In fact, a cobordism trick using [CE20, Laz20b] allows us to reduce the connected sum result to the case where the contact manifold we begin with is Weinstein fillable. In spite of this, the full proof does require that symplectic cohomology can be defined for general strong fillings, which has now been established using for example Pardon’s VFC package [Par19].
The proof of Theorem C yielding obstructions to Weinstein fillings are technically simpler, as one has exact fillings by assumption, and one can follow the uniqueness proofs without significant changes. Here the key step is to ensure that the surgery procedure can be applied to reach a standard sphere in the first place.
@@
In Section 3, we describe then the notion of algebraic tightness, together with relevant properties, and prove that the Bourgeois construction is a rich source of contact manifolds satisfying this property.
Section 4 contains results on obstructions to strong fillability via symplectic cohomology and Section 5 contains the proof of Theorem B and Theorems A and D. Lastly, Section 6 contains the proofs of Theorems C and E.
@@ Acknowledgements
A. Moreno received support from the National Science Foundation under Grant No. DMS-1926686, and is currently supported by the Sonderforschungsbereich TRR 191 Symplectic Structures in Geometry, Algebra and Dynamics, funded by the DFG (Projektnummer 281071066 – TRR 191), and also by the DFG under Germany’s Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster). Z. Zhou is supported by National Natural Science Foundation of China under Grant No. 12288201 and 12231010. F. Gironella gratefully acknowledges the support of the Sonderforschungsbereich Higher Invariants 1085. This project has also received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 772479), as well as from the Agence Nationale de la Recherche (ANR) under the Project COSY (ANR-21-CE40-0002-04). The authors are moreover grateful to Peter Albers and Yasha Eliashberg, for productive discussions and interest in the project, and to Frédéric Bourgeois, Roger Casals, Otto van Koert, Oleg Lazarev, Alex Oancea, for helpful feedback on an earlier version of this paper.
@@ 2. Bourgeois contact manifolds
In this section, we recall the Bourgeois construction [Bou02b], following the presentation that was given in [BGM22].
Consider a closed, oriented, connected smooth manifold $M^{2n-1}$ and an open book decomposition that we denote $(B,\theta)$, together with a defining map $\Phi\colon\thinspace M\rightarrow\mathbb{R}^{2}$ so that each $z\in\operatorname{int}(\mathbb{D}^{2})$ is a regular value. Here, $B\subset M$ is a closed codimension-2 submanifold, $\theta\colon M\backslash B\rightarrow\mathbb{S}^{1}$ is a fiber bundle, and $\Phi$ is such that $\Phi^{-1}(0)=B$ and $\theta=\Phi/\left|\Phi\right|$. Let us also denote by $\rho$ the norm $|\Phi|$.
@@
The contact form $\alpha_{BO}$ on $M\times\mathbb{T}^{2}$ will be called the Bourgeois form associated to $(\alpha,\Phi)$ in what follows. It will be useful to also think in terms of abstract open books, i.e. in terms of the Liouville page $\Sigma=(\Sigma,d\lambda)$ and the compactly supported symplectic monodromy $\psi$. In this case we write $OB(\Sigma,\psi)$ for the contact manifold obtained via the Thurston-Winkelnkemper construction, and we denote $BO(\Sigma,\psi)=(M\times\mathbb{T}^{2},\ker\alpha_{BO})$. We invite the reader to consult [LMN19] for details on the correspondence between Bourgeois contact structures for geometric and abstract open books.
@@
where $\Omega$ is an area form on $\mathbb{T}^{2}$. This can be seen via an explicit homotopy; see for instance [Gir20a, Lemma 4.1.(1)]. In particular if $c_{1}(\xi)$ is torsion, then the same is true for the Bourgeois contact structure, independently of the auxiliary choice of open book.
@@ 2.1. Reeb dynamics of Bourgeois contact structures
In this somewhat technical section, we consider the Reeb dynamics of the Bourgeois contact forms more closely. First of all applying a lemma of Giroux (see e.g. [DGZ14, Section 3]), on a neighborhood of the form $\mathcal{N}=B\times\mathbb{D}^{2}\times\mathbb{T}^{2}$ of $B$ where $\rho=|\Phi|$ coincides with the radial coordinate of $\mathbb{D}^{2}$, the Bourgeois contact form looks like
$\alpha_{BO}$ $=\alpha+\rho(r)(\cos(\theta)dq_{1}-\sin(\theta)dq_{2})$
$=h_{1}(r)\alpha_{B}+h_{2}(r)d\theta+r(\cos(\theta)dq_{1}-\sin(\theta)dq_{2}).$
@@
We also point out that there is a natural (orientation preserving) diffeomorphism
@@
We then have the following global description of the Reeb vector field (c.f. [Bou02b, Gir20a]):
@@
The Reeb vector field of the contact form $\alpha_{BO}$ is given by
$R_{BO}=\mu(r)R_{B}+\nu(r)[\sin(\theta)\partial_{q_{1}}-\cos(\theta)\partial_{q_{2}}],$ (2)
@@
The computation in [Bou02b] shows that, in order for $\alpha_{BO}$ to be a contact form in the neighborhood $\mathcal{N}\simeq B\times D^{*}\mathbb{T}^{2}$ of $B\times\mathbb{T}^{2}$ where $\alpha=h_{1}\alpha_{B}+h_{2}d\theta$, it is in fact enough that $h_{1}(h_{1}-h_{1}^{\prime})>0$, a condition which only depends on $h_{1}$. In particular, one can homotope the pair $(h_{1},h_{2})$ in a compactly supported way among pairs of functions satisfying this condition, and this will result in a homotopy of contact forms (hence isotopy by Gray’s stability) on $M\times\mathbb{T}^{2}$, independently of the fact that the resulting $\alpha$ on $M$ might not be adapted to the open book or even a contact form. Moreover, the explicit formula in Lemma 2.3 still holds for the homotoped contact form, as the explicit computation does not use any specific property of the pair of functions $(h_{1},h_{2})$ listed above. In particular, up to homotopy one can achieve the following form, which will be useful below:
Lemma 2.4.
For $\delta>0$ sufficiently small, up to a deformation among contact structures on $M\times\mathbb{T}^{2}$ supported in the neighbourhood of radius $2\delta$ of $B\times\mathbb{T}^{2}$, one can assume that
  • $h_{1}(r)=1$ for $r\leq\delta$,
  • $h_{1}(h_{1}-h_{1}^{\prime})>0$ everywhere,
  • $h_{2}(r)=0$ for $r\leq 3\delta/2$.
In particular, under the diffeomorphism in Equation 1, $\alpha_{BO}$ coincides with $\alpha_{B}+\lambda_{std}$ for $r=\sqrt{p_{1}^{2}+p_{2}^{2}}\leq\delta$, where $\lambda_{std}=p_{1}dq_{1}+p_{2}dq_{2}$ is the standard Liouville form on $D^{*}\mathbb{T}^{2}$. Moreover, Equation 2 also holds for the Reeb vector field of the deformed contact form; in particular, it coincides with $R_{B}$ for $r\leq\delta$.

### 14 Mar 2026 v5

@@ 1. Introduction
In contact topology, there is an important dichotomy dividing contact structures into those that are overtwisted and those that are tight. According to [Eli89, BEM15], the former are in fact topological and not geometric objects, in the sense that they satisfy $h$-principles so that their existence and classification can be reduced to classical obstruction theory and algebraic topology. On the other hand, tight contact structures are inherently geometric, hence in general much harder to construct, and let alone classify. There is, however, an important sufficient condition for tightness: every contact manifold which is symplectically fillable is necessarily tight [Gro85, Eli90, Nie06, BEM15]. Here we consider strong fillings, i.e. symplectic manifolds such that the contact structure at the boundary is defined by the contraction of the symplectic form with a Liouville vector field, locally defined near the boundary. A natural question is then to understand the difference between the classes of tight and strongly fillable contact structures. Our first result addresses this question in broad generality, on manifolds of dimension at least five.
@@
If $(M^{2n+1},\xi)$ is tight with $n\geq 3$, then $M$ admits a tight, non-strongly-fillable contact structure in the same almost contact class. If $n=2$, the same holds, under the additional assumption that the first Chern class of $\xi$ is zero.
The assumption on the dimension in Theorem A is necessary, since in dimension three the statement is false. For example, the sphere $\mathbb{S}^{3}$ admits a unique tight contact structure [Eli92], which is also symplectically fillable by the standard Darboux ball. Historically the first examples of tight contact structures that are not symplectically fillable were found in dimension $3$ by Entyre-Honda [EH02] and on a special class of product manifolds in dimensions 5 and greater by Massot-Niederkrüger-Wendl [MNW13].
@@
For every $n\geq 2$, the sphere $\mathbb{S}^{2n+1}$ admits a tight, non-strongly-fillable contact structure that is homotopically standard.
Here, homotopically standard means homotopic to the standard contact structure among almost contact structures (the algebraic topological object underlying contact structures, in terms of which the $h$-principles in [Eli89, BEM15] are formulated). We then deduce Theorem A by considering connected sums with these non-fillable examples on spheres to preserve the smooth topology but alter the properties of the contact structure. Of course, this requires some significant effort since both the obstruction to fillability as well as tightness do not behave well under connected sums a priori.
@@
The notion of fillability comes with several different nuances, namely, one has notions of Weinstein fillability, Liouville fillability and strong fillability. Thus ordered, each notion is stronger than the following one, and all of them in fact imply tightness. One thus has the following natural hierarchy of contact manifolds:
$\{\mbox{Weinstein fillable}\}\subseteq\{\mbox{Liouville fillable}\}\subseteq\{\mbox{strongly fillable}\}\subseteq\{\mbox{tight}\}.$
There has been a considerable amount of work over almost 30 years in studying these inclusions, and they are known to be strict in any odd dimension $\geq 3$ [Eli96, EH02, Ghi05, Gay06, NW11, Bow12, MNW13, BCS14, Zho21a, BGM22, GNE22]. On the other hand, all previously known examples of contact structures with exotic fillability properties come from special geometric constructions, all of which again require the presence of non-trivial topology in the underlying smooth manifold. Now, as a consequence of Theorem A and Theorem B, the final inclusion is strict for all manifold of dimension $\geq 7$, as soon as this admits at least one tight contact structure.
@@
In fact, there is also the further notion of weak fillability, which, as the name suggests, is the weakest possible notion of fillability. Since this in fact agrees with strong fillability on spheres due to [MNW13, Proposition 6], we do not consider weak fillability here. In particular, even though strictly speaking in (the proof of) Theorem B we obstruct strong fillability, this equivalence of strong and weak fillability for spheres implies that those contact spheres in fact are also not weakly fillable.
@@
By taking connected sum with examples on spheres, we can further deduce that there are abundant Liouville fillable contact manifolds without Weinstein fillings in the case that the first Chern class is torsion (cf. Theorem E below). Thus, other than addressing the $5$-dimensional case, it remains to determine the existence of strongly but not Liouville fillable contact structures on $\mathbb{S}^{2n+1}$, for all $n\geq 2$.
@@
After one constructs examples of contact structures with exotic fillability properties, it is natural to try to produce infinitely many such examples as well. In this direction, by pushing the techniques used for the proofs of the previous statements further, namely by taking contact connected sums with some other judiciously chosen Weinstein fillable contact structures on the sphere, we obtain the following.
@@
Let $n\geq 5$, and suppose that $M^{2n+1}$ admits a Weinstein fillable contact structure with torsion first Chern class. Then, $M$ admits infinitely many non-isomorphic, tight, and not strongly fillable contact structures in the same almost contact class.
In particular, spheres of dimension at least $11$ admit infinitely many tight non-fillable contact structures which are homotopically standard. One also has a similar result for Liouville, but not Weinstein fillable contact structures.
@@
To prove Theorem B, i.e. to obtain a tight and non-fillable contact structure on $\mathbb{S}^{2n+1}$, we give a geometric construction, as follows. We consider the Bourgeois construction on a sphere $\mathbb{S}^{2n-1}$ (with respect to a suitable open book), which yields a contact structure on $\mathbb{S}^{2n-1}\times\mathbb{T}^{2}$, and then we kill the topology coming from the $\mathbb{T}^{2}$-factor via isotropic and flexible surgeries to obtain a contact $2n+1$ dimensional sphere. The key point, which we now discuss briefly, is to ensure that tightness and non-fillability of the starting Bourgeois contact manifold are preserved under the contact surgeries.
Tightness. In general, tightness is known to be preserved under surgery only in dimension $3$ by a result of Wand [Wan15]. Here, we circumvent the lack of an analogous result in higher dimensions by using the stronger notion of algebraic tightness, i.e. non-vanishing of contact homology, and the fact that this is known to be preserved under contact surgeries and to imply tightness (see Proposition 3.2). Therefore, any contact manifold obtained via contact surgery from an algebraically tight one is in particular (algebraically) tight.
In our situation, it hence suffices to arrange that the starting Bourgeois contact structure is algebraically tight, which can be done as follows. When $n=2$, we follow the same strategy in [BGM22], leveraging the existence of a certain exact cobordism with hypertight concave end and convex end, the desired Bourgeois manifold, and appealing to the functoriality of contact homology. When $n\geq 3$, we can make the Bourgeois construction satisfy certain index-positivity condition (more precisely, one can arrange them to be $1$-asymptotically dynamically convex, as introduced in [Laz20a, Zho21b]; see Lemma 2.8), which implies algebraic tightness.
Non-fillability. The obstruction to fillability is derived from restrictions of fillings for certain flexibly Weinstein fillable contact manifolds. Taking the celebrated Eliashberg–Floer–McDuff theorem [McD91] as a starting point, many contact manifolds are expected to have unique Liouville/symplectically aspherical fillings, at least up to diffeomorphism. Up until now, results in this direction mainly apply to the contact boundary of a split manifold of the form $V\times\mathbb{D}^{2}$, as well as those obtained via subcritical/flexible surgeries; cf. [OV12, BGZ19, Zho21b, Zho23b, BGM22, GKZ23]. If one considers the more general class of strong fillings, these are no longer unique, as one can take a blow-up of any given filling without affecting the boundary.
However, it is plausible that all the ambiguity of fillings of $\partial(V\times\mathbb{D}^{2})$ only comes from birational surgeries like blow-ups, which will increase the complexity of the topology of the filling; see [Zho21b, §8] for supporting evidence. Moreover, Eliashberg [Eli90] showed in dimension $3$, that if $W$ is a symplectic filling of $Y\#Y^{\prime}$, then $W$ is obtained from attaching a $1$-handle to a symplectic filling $W^{\prime}$ of $Y\sqcup Y^{\prime}$ (which might be connected, cf. [Bow12]). Although the higher-dimensional analogue has not been established, partial generalizations were obtained by Ghiggini-Niederkrüger-Wendl [GNW16]. Incarnations of these expectations at the level of homology lead to Theorem 4.1, putting homology constraints on putative strong fillings, which serve as our obstructions to the existence of strong fillings.
@@
In [BGM22], the first three authors used moduli spaces of closed spheres that appear after applying a certain capping procedure, similar as to what is done in [MNW13] to obstruct fillability of certain Bourgeois contact manifolds. In this paper, we instead use a partial cap (described in Proposition 2.13) to build a strong cobordism from the Bourgeois contact manifold as well as the exotic sphere to a flexibly fillable manifold, for which we can apply Theorem 4.1. The full proof does require that symplectic cohomology can be defined for general strong fillings, which has now been established, for instance, using Pardon’s VFC package [Par19].
For the proof of Theorem A, one simply takes the original tight contact manifold, and performs a connected sum with the exotic sphere of Theorem B. The subtle point is to ensure that tightness and non-fillability are preserved under this connected sum. The proof of non-fillability uses the fact that the homological obstructions can be made to persist under connected sum, as explained above. Tightness is again reduced to algebraic tightness: one can assume without loss of generality that the original contact manifold is strongly fillable (otherwise there is nothing to prove), and therefore it is algebraically tight; then, we can appeal to the fact (see Proposition 3.2) that algebraic tightness is preserved under connected sum, that leads in particular to tightness of the considered connected sum, as desired.
The proof of Theorem C yielding obstructions to Weinstein fillings is technically simpler. The construction is the following: we start from a Liouville domain of the form $V^{2n}=N\times[-1,1]$ as constructed by [MNW13]. We let $M=\partial(V\times\mathbb{D}^{2})$, which admits a flexible Weinstein cobordism to a homotopy sphere by Lemma 6.1. Moreover, up to taking finitely many self-connected sums, it can be assumed to be the standard smooth sphere. The cohomology of the filling resulting from stacking the cobordism on top of $V\times\mathbb{D}^{2}$ is non-trivial in degree 2n12n-1, which evaluates non-trivially on the fundamental class of $N$. By results in [Zho21b], the cohomology of any Weinstein filling of this contact sphere is the same as the cohomology of the natural Liouville one, so in particular non-trivial in degree 2n1>n+12n-1>n+1 (as $n\geq 3$), which obstructs the Weinstein property. Therefore, we obtain a Liouville but non-Weinstein fillable contact structure. The standard almost contact structure is also achieved by a connected sum trick.
Finally, the examples in Theorem D are also obtained by considering connected sums of the examples from Theorem A with some flexibly filled examples that are constructed in a somewhat ad-hoc manner. Tightness and non-fillability follow from the same line as for Theorem A, and the tool used to distinguish them up to contactomorphism is positive symplectic cohomology, which is a contact invariant in this setting thanks to $1$-asymptotic dynamical convexity and can be defined without using a Liouville filling by [CO18]. Similarly, Theorem E follows from connect-summing the examples of Theorem C with infinitely many homotopically trivial flexibly fillable contact spheres from [Laz20a], and they are distinguished using again positive symplectic cohomology but this time of their natural Liouville fillings, that is an invariant of the contact structure by [Laz20a].
@@
In Section 3, we describe the notion of algebraic tightness, together with relevant properties, and prove that the Bourgeois construction is a rich source of contact manifolds satisfying this property.
Section 4.1 contains results on obstructions to strong fillability via symplectic cohomology and Section 5 contains the proof of Theorem B and Theorems A and D. Lastly, Section 6 contains the proofs of Theorems C and E.
@@ Acknowledgements
The authors are grateful to Peter Albers and Yasha Eliashberg, for productive discussions and interest in the project, to Frédéric Bourgeois, Roger Casals, Otto van Koert, Oleg Lazarev, Alex Oancea, for helpful feedback on an earlier version of this paper, and to anonymous referees for many suggestions that improved this paper.
J. Bowden gratefully acknowledges the support from the Heisenberg-Grant (BO 4423/4-1) of the German Research Council. F. Gironella is supported by the region Pays de la Loire, via the project Étoile Montante 2023 SymFol, and gratefully acknowledges the support of the Sonderforschungsbereich Higher Invariants 1085. A. Moreno received support from the National Science Foundation under Grant No. DMS-1926686, and is currently supported by the Sonderforschungsbereich TRR 191 Symplectic Structures in Geometry, Algebra and Dynamics, funded by the DFG (Projektnummer 281071066 – TRR 191), and also by the DFG under Germany’s Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster). Z. Zhou is supported by National Key R&D Program of China under Grant No.2023YFA1010500 and National Natural Science Foundation of China under Grant No. 12288201 and 12231010. This project has also received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 772479), as well as from the Agence Nationale de la Recherche (ANR) under the Project COSY (ANR-21-CE40-0002-04).
@@ 2. Bourgeois contact manifolds
In this section, we recall the Bourgeois construction [Bou02a], following the presentation that was given in [BGM22].
Consider a closed, oriented, connected smooth manifold $M^{2n-1}$ and an open book decomposition that we denote $(B,\theta)$, together with a defining map $\Phi\colon\thinspace M\rightarrow\mathbb{R}^{2}$ such that each $z\in\operatorname{int}(\mathbb{D}^{2})$ is a regular value. Here, $B\subset M$ is a closed codimension-2 submanifold, $\theta\colon M\backslash B\rightarrow\mathbb{S}^{1}$ is a fiber bundle, and $\Phi$ is such that $\Phi^{-1}(0)=B$ and $\theta=\Phi/\left|\Phi\right|$. Let us also denote by $\rho$ the norm $|\Phi|$.
@@
The contact form $\alpha_{BO}$ on $M\times\mathbb{T}^{2}$ will be called the Bourgeois form associated to $(\alpha,\Phi)$ in what follows. It will be useful to also think in terms of abstract open books, i.e. in terms of the Liouville page $\Sigma=(\Sigma,d\lambda)$ and the compactly supported symplectic monodromy $\psi$. In this case, we write $OB(\Sigma,\psi)$ for the contact manifold obtained via the Thurston-Winkelnkemper construction, and we denote $BO(\Sigma,\psi)=(M\times\mathbb{T}^{2},\ker\alpha_{BO})$. We invite the reader to consult [LMN19] for details on the correspondence between Bourgeois contact structures for geometric and abstract open books.
@@
where $\Omega$ is an area form on $\mathbb{T}^{2}$. This can be seen via an explicit homotopy; see, for instance [Gir20a, Lemma 4.1.(1)]. In particular, if $c_{1}(\xi)$ is torsion, then the same is true for the Bourgeois contact structure, independently of the auxiliary choice of open book.
@@ 2.1. Reeb dynamics of Bourgeois contact structures
In this section, we consider the Reeb dynamics of the Bourgeois contact forms more closely. First of all applying a lemma of Giroux (see e.g. [DGZ14, Section 3]), on a neighborhood of the form $\mathcal{N}=B\times\mathbb{D}^{2}\times\mathbb{T}^{2}$ of $B$ where $\rho=|\Phi|$ coincides with the radial coordinate of $\mathbb{D}^{2}$, the Bourgeois contact form looks like
$\alpha_{BO}$ $=\alpha+\rho(r)(\cos(\theta){\mathrm{d}}q_{1}-\sin(\theta){\mathrm{d}}q_{2})$
$=h_{1}(r)\alpha_{B}+h_{2}(r){\mathrm{d}}\theta+r(\cos(\theta){\mathrm{d}}q_{1}-\sin(\theta){\mathrm{d}}q_{2}).$
@@
We also point out that there is a natural (orientation-preserving) diffeomorphism
@@
We have the following global description of the Reeb vector field (c.f. [Bou02a, Gir20a]) on a Bourgeois manifold.
@@
The Reeb vector field of the contact form $\alpha_{BO}$ is given globally by
$R_{BO}=\mu(r)R_{B}+\nu(r)[\cos(\theta)\partial_{q_{1}}-\sin(\theta)\partial_{q_{2}}],$ (2)